# Self-Supervised Transformers as Iterative Solution Improvers for Constraint Satisfaction

### v1

@@ Abstract
We present a Transformer-based framework for Constraint Satisfaction Problems (CSPs). CSPs find use in many applications and thus accelerating their solution with machine learning is of wide interest. Most existing approaches rely on supervised learning from feasible solutions or reinforcement learning, paradigms that require either feasible solutions to these NP-Complete CSPs or large training budgets and a complex expert-designed reward signal. To address these challenges, we propose ConsFormer, a self-supervised framework that leverages a Transformer as a solution refiner. ConsFormer constructs a solution to a CSP iteratively in a process that mimics local search. Instead of using feasible solutions as labeled data, we devise differentiable approximations to the discrete constraints of a CSP to guide model training. Our model is trained to improve random assignments for a single step but is deployed iteratively at test time, circumventing the bottlenecks of supervised and reinforcement learning. Our method can tackle out-of-distribution CSPs simply through additional iterations.
@@ 1 Introduction
To address these challenges, we introduce ConsFormer, a Transformer-based self-supervised framework for solving CSPs. ConsFormer learns to iteratively refine variable assignments through a self-supervised training paradigm that approximates discrete constraints with continuous differentiable penalty functions. Our model is trained to improve an initial random assignment in a single refinement step, but is applied iteratively at test time. While a single refinement step may not yield a feasible solution, a sufficiently large number of improving iterations (on average) does. Examples of ConsFormer solutions are shown in Figure 1.
Transformers provide a natural fit for this approach due to their strong generalization capabilities and their ability to process structured data efficiently (Lewkowycz et al., 2022; Achiam et al., 2023). They are particularly effective at learning with tokenized inputs, making them well-suited for combinatorial problems formulated in the Constraint Programming paradigm (Rossi et al., 2006). Furthermore, recurrence has been shown to enhance the generalization abilities of Transformers (McLeish et al., 2024; Fan et al., 2024), reinforcing their suitability for our setting.
Our iterative solution improvement strategy enables ConsFormer to generalize beyond its training distribution, effectively solving out-of-distribution (OOD) CSP instances simply by performing more refinement steps. We evaluate ConsFormer on a diverse set of CSP problems, including Sudoku, Graph Coloring, and Nurse Scheduling, demonstrating its ability to generalize across problem domains.
The following high-level findings summarize our contributions:
@@ 2.2 Related Work
A common recipe for Reinforcement Learning (RL) in constraint solving is to express the problem with a graph which is then processed using Graph Neural Networks (GNN). The GNN’s weights are updated using RL based on a reward function expressing an objective function and/or constraint satisfaction (Dai et al., 2017; Chalumeau et al., 2021; Li et al., 2024b; Boisvert et al., 2024). For example, Tönshoff et al. (2023) converts a CSP instance into a tripartite variable-domain-constraint graph which is then solved using a GNN trained by RL. Wu et al. (2022) use a Transformer architecture to learn discrete transformation steps with RL for routing problems. However, RL approaches require significant computational resources for training, as well as an expertly designed reward signal for each problem.
Non-RL based methods require addressing the non-differentiability of discrete constraints. Yang et al. (2022) use the straight-through estimator (Bengio et al., 2013) for logical constraints and Tang et al. (2024) explore a similar approach for mixed-integer non-linear programs. Toenshoff et al. (2021) devise a continuous relaxation for binary constraints (i.e., constraints involving two variables) which are used to guide a recurrent GNN to generate solutions; this approach is limited in applicability as many CSPs of interest have non-binary constraints. Bai et al. (2021) design continuous relaxations for some constraint classes in conjunction with a reconstruction loss to tackle a visual Sudoku problem; it is unclear how their architecture can be adapted to CSP solving in general.
@@
Continuous relaxations have been used effectively to approximate discrete functions. For example, T-norm has been widely implemented as a continuous approximation for discrete binary logic operations (Petersen et al., 2022; Giannini et al., 2023; Gimelfarb et al., 2024). Petersen et al. (2021) introduced continuous relaxations for discrete algorithms, such as if-statements and while-loops. Similarly, Chen et al. (2019) proposed entropy-based relaxations to handle discrete constraints.
@@
The incorporation of recurrency has been shown to improve a model’s generalization. Bansal et al. (2022) implement recurrent ResNet blocks to solve simple logic puzzles and show that increasing recurrent steps at test-time allows generalization to harder unseen tasks. Recurrency was introduced to the Transformer architecture by sharing weights across Transformer layers (Dehghani et al., 2019; Takase & Kiyono, 2023), yielding improved generalization capabilities on arithmetic and logic-based string manipulation tasks (McLeish et al., 2024; Fan et al., 2024). Our method differs from the existing work as recurrency is only introduced during test-time deployment.
@@ 3 ConsFormer: a Single-Step Self-Supervised Transformer
Figure 2: ConsFormer: a Single-Step Self-Supervised Transformer for CSP. A CSP instance is transformed into a set of input embeddings composed of the variable value encoding, index-based Absolute Position Encoding, and a selected update set embedding indicating the variables to be updated. The input is processed by $L$ Transformer layers, incorporating the binary constraint graph as Relative Positional Encodings for attention. The output values are used to compute a differentiable self-supervised loss on constraint violation. Although trained to perform one step of solution improvement, ConsFormer can be deployed iteratively at test time, improving the odds of solving instances never seen during training.
@@ 3.1 Input Representation
Variable assignments as tokens.
Let $X=\{x_{1},x_{2},\dots,x_{n}\}$ be the set of variables in a CSP, each of which has a finite domain $D_{i}$. A variable assignment $x_{i}=v$, $v\in D_{i}$, is treated as a token. A learnable embedding $\mathbf{e}(v)$ is assigned to each unique value $v\in\bigcup_{i=1}^{n}D_{i}$. The input variable assignment, represented as $\mathbf{X}=\{x_{1}=v_{1},x_{2}=v_{2},\dots,x_{n}=v_{n}\}$, forms the input token set to the Transformer. Thus, the input embedding set is given by
@@
Representing variable indices with Absolute Positional Encoding.
Transformers use Absolute Positional Encoding (APE) to represent the position of tokens in a sequence. For CSPs, we use APE to encode the indices of variables. If the indices of a variable $x_{i}$ are multi-dimensional, we concatenate the positional encodings for each dimension. Specifically, let $x_{i_{1},i_{2},\dots,i_{k}}$ denote a variable with $k$-dimensional indices $(i_{1},i_{2},\dots,i_{k})$. The APE for this variable is computed as:
@@
where $\text{PE}(i_{k})$ is the positional encoding for dimension $k$. For example, in Sudoku, a variable $x_{12}$ would have an APE formed by concatenating the encodings for row 1 and column 2. This approach is inspired by the APE design in Vision Transformers (Carion et al., 2020; Li et al., 2024a).
Constraint relations as Relative Positional Encoding.
A Relative Positional Encoding (RPE) is typically used by Transformers to capture the positional relationship between tokens independently of their absolute positions in a sequence. For CSPs, we use RPE to encode the constraint relationships between variables. Specifically, we represent the CSP constraints as a binary constraint graph $\mathbf{G}=(V,E)$, where $V=\{1,2,\dots,n\}$ corresponds to the variables and $E$ contains edges between pairs of variables that participate together in at least one constraint of the CSP.
@@ 3.2 A Single-Step Transformer Architecture
where $\mathbf{e}(v_{i})$ is the token embedding of variable $x_{i}$’s current value $v_{i}$ as described in Equation 1, and $\text{APE}(x_{i})$ is its positional encoding as computed in Equation 2. $\alpha,\beta,\gamma$ are learnable scalars that allow the model to balance the contributions of each encoding, inspired by Li et al. (2024a). The set of embeddings for all variables forms the input set $\mathbf{H}^{(0)}=\{\mathbf{h}_{1}^{(0)},\mathbf{h}_{2}^{(0)},\dots,\mathbf{h}_{n}^{(0)}\}$.
@@
ConsFormer employs a self-attention mechanism to compute updated representations of variables based on other variables. For each variable token $\mathbf{h}_{i}^{(l)}\in\mathbb{R}^{h\times 1}$ at layer $l$, the self-attention mechanism proceeds as follows:
  1. 1.
@@ 3.3 Self-supervised Loss Function
where hyperparameter $\lambda_{i}$ is the weight assigned to $p_{i}$, and $f$ is an optional operation to transform the penalty for better learning. In practice, we implemented the common quadratic penalty, $f(x)=x^{2}$. The discrete constraints and their relaxed continuous counterparts we implemented for our experiments are shown in Table 1, and numerical examples of valid and invalid assignments for each constraint can be found in Appendix A.
@@ 3.4 Iterative Test-Time Deployment
In this sense, ConsFormer can be viewed as performing a single step of neural local search to improve the candidate solution. Our experiments focus on basic iterative solution refinement in one continuous sequence, without additional augmentations. However, this capability can be combined with techniques such as backtracking and random restarts (Rossi et al., 2006) to create a neuro-symbolic solver. Another possible extension is to incorporate ConsFormer as an evolutionary algorithm (Holland, 1992) utilizing the Transformer’s parallel processing ability to update a pool of candidate solutions all at once. We leave these explorations as future work.
@@ 4.1 Problems
Graph Coloring is the problem of finding an assignment of colors to vertices in a graph such that no two neighboring nodes share the same color. The problem is defined by the graph’s structure and the number of available colors $k$. We generate two sets of graph instances for $k=5$ and $k=10$ following a similar procedure as (Tönshoff et al., 2023) (See Appendix C for details). Training graphs have 50 vertices for $k=5$ and 100 vertices for $k=10$ whereas OOD graphs have 100 for $k=5$ and 200 for $k=10$. We use inequality constraints of the form $x_{i}\neq x_{j}$ for an edge between nodes $i$ and $j$ and their penalty in Table 1 to represent the coloring constraints.
@@
Table 2: Performance comparison for Sudoku. In-distribution test instances contain 1K instances from the SATNet dataset, OOD refers to Out-of-Distribution evaluation on the RRN test dataset which contains 18K instances. *Values reported in (Du et al., 2024).
@@ 4.3 Results
Sudoku.
Table 2 reports the performance of ConsFormer and various neural methods on the Sudoku task. ConsFormer solves 100% of the Sudoku tasks from the in-distribution test dataset. On the harder out-of-distribution dataset, ConsFormer significantly outperforms all learned methods, demonstrating superior generalization capabilities. ConsFormer achieves instance solve rates of 65.88% and 77.74% with 2k and 10k iterations, respectively.
@@
Graph Coloring.
Table 3 summarizes the performance of OR-Tools, ANYCSP (Tönshoff et al., 2023), and ConsFormer on Graph Coloring instances. OR-Tools is a state-of-the-art traditional solver for constraint programming applications and serves as a strong baseline (Perron & Didier, ), achieving 100% on Sudoku instances. We ran test instances sequentially through all models with a 10-second timeout, though ConsFormer can process instances significantly faster if processed in batches due to the Transformer architecture. We note that the harder dataset is not out of distribution for OR-Tools, since it solves each task individually and is not a learning-based solver.
@@
On the more challenging Graph Coloring with $k=10$ datasets, ConsFormer surpasses OR-Tools in performance, showcasing the advantage of learned heuristic approaches: it may not surpass state-of-the-art solvers on smaller instances, but it excels in complex cases under short time limits—crucial for real-world applications. Surprisingly, ANYCSP failed to solve any instances within 10 seconds for both datasets, underscoring the scalability limitations of its graph-based representation and the difficulty of training with RL.

### v2

@@ Abstract
We present a Transformer-based framework for Constraint Satisfaction Problems (CSPs). CSPs find use in many applications and thus accelerating their solution with machine learning is of wide interest. Most existing approaches rely on supervised learning from feasible solutions or reinforcement learning, paradigms that require either feasible solutions to these NP-Complete CSPs or large training budgets and a complex expert-designed reward signal. To address these challenges, we propose ConsFormer, a self-supervised framework that leverages a Transformer as a solution refiner. ConsFormer constructs a solution to a CSP iteratively in a process that mimics local search. Instead of using feasible solutions as labeled data, we devise differentiable approximations to the discrete constraints of a CSP to guide model training. Our model is trained to improve random assignments for a single step but is deployed iteratively at test time, circumventing the bottlenecks of supervised and reinforcement learning. Experiments on Sudoku, Graph Coloring, Nurse Rostering, and MAXCUT demonstrate that our method can tackle out-of-distribution CSPs simply through additional iterations.
@@ 1 Introduction
To address these challenges, we introduce ConsFormer, a Transformer-based self-supervised framework for solving CSPs. Inspired by Constraint-based Local Search (Hentenryck & Michel, 2009), ConsFormer learns to iteratively refine variable assignments through a self-supervised training paradigm that approximates discrete constraints with continuous differentiable penalty functions. Our model is trained to improve an initial random assignment in a single refinement step, but is applied iteratively at test time. While a single step may not yield a feasible solution, a sufficiently large number of improving iterations (on average) does. Examples of ConsFormer solutions are shown in Figure 1.
Transformers provide a natural fit for this approach due to their strong generalization capabilities and their ability to process structured data efficiently (Lewkowycz et al., 2022; Achiam et al., 2023). They are particularly effective at learning with tokenized inputs, making them well-suited for combinatorial problems formulated in the Constraint Programming paradigm (Rossi et al., 2006). Furthermore, recurrence has been shown to enhance the generalization abilities of Transformers (McLeish et al., 2024; Fan et al., 2025), reinforcing their suitability for our setting.
Our iterative solution improvement strategy enables ConsFormer to generalize beyond its training distribution, effectively solving out-of-distribution (OOD) CSP instances simply by performing more refinement steps. We evaluate ConsFormer on a diverse set of CSP problems, including Sudoku, Graph Coloring, and Nurse Scheduling, demonstrating its ability to generalize across problem domains. Our implementation is available on GitHub [^1] https://github.com/khalil-research/ConsFormer.
The following high-level findings summarize our work:
@@ 2.2 Related Work
A common recipe for Reinforcement Learning (RL) in constraint solving is to express the problem with a graph which is then processed using Graph Neural Networks (GNN). The GNN’s weights are updated using RL based on a reward function expressing an objective function and/or constraint satisfaction (Dai et al., 2017; Chalumeau et al., 2021; Li et al., 2024; Boisvert et al., 2024). For example, Tönshoff et al. (2023) converts a CSP instance into a tripartite variable-domain-constraint graph which is then solved using a GNN trained by RL. Similarly, Yolcu & Póczos (2019) represent SAT problems using a clause-variable graph. Wu et al. (2022) use a Transformer architecture to learn discrete transformation steps with RL for routing problems. However, RL approaches require significant computational resources for training, as well as an expertly designed reward signal for each problem.
Non-RL based methods require addressing the non-differentiability of discrete constraints. Yang et al. (2022) use the straight-through estimator (Bengio et al., 2013) for logical constraints and Tang et al. (2024) explore a similar approach for mixed-integer non-linear programs. Toenshoff et al. (2021) devise a continuous relaxation for binary constraints (i.e., constraints involving two variables) which are used to guide a recurrent GNN to generate solutions; this approach is limited in applicability as many CSPs of interest have non-binary constraints. Bai et al. (2021) design continuous relaxations for some constraint classes in conjunction with a reconstruction loss to tackle a visual Sudoku problem; it is unclear how their architecture can be adapted to CSP solving in general. Self-supervised learning has been successfully applied in continuous domains (Donti et al., ; Park & Van Hentenryck, 2023), as well as for SAT (Ozolins et al., 2022).
@@
Continuous relaxations have been used effectively to approximate discrete functions. For example, T-norm has been widely implemented as a continuous approximation for discrete binary logic operations (Petersen et al., 2022; Giannini et al., 2023; Gimelfarb et al., 2024). Petersen et al. (2021) introduced continuous relaxations for discrete algorithms, such as if-statements and while-loops. Combinatorial constraints have also been approximated using entropy-based relaxations (Chen et al., 2019), probabilistic methods (Karalias & Loukas, 2020; Bu et al., 2024), and set function extensions (Karalias et al., 2022).
@@
The incorporation of recurrency has been shown to improve a model’s generalization. Bansal et al. (2022) implement recurrent ResNet blocks to solve simple logic puzzles and show that increasing recurrent steps at test-time allows generalization to harder unseen tasks. Recurrency was introduced to the Transformer architecture by sharing weights across Transformer layers (Dehghani et al., 2019; Takase & Kiyono, 2023), yielding improved generalization capabilities on arithmetic and logic-based string manipulation tasks (McLeish et al., 2024; Fan et al., 2025). Our method differs from the existing work as recurrency is only introduced during test-time deployment.
@@ 3 ConsFormer: a Single-Step Self-Supervised Transformer
Figure 2: ConsFormer: a Single-Step Self-Supervised Transformer for CSP. A CSP instance is transformed into a set of input embeddings composed of the variable assignment value encoding, index-based Absolute Position Encoding, and a selected update set embedding $e_{s}$ indicating the variables to be updated. The input is processed by $L$ Transformer layers, incorporating the binary constraint graph as Relative Positional Encodings for attention. The output values are used to update the variable assignments which are then used to compute a differentiable self-supervised loss on constraint violation. Although trained to perform one step of solution improvement, ConsFormer can be deployed iteratively at test time, improving the odds of solving instances never seen during training.
@@ 3.1 Input Representation
Variable assignments as tokens. Let $X=\{x_{1},x_{2},\dots,x_{n}\}$ be the set of variables in a CSP, each of which has a finite domain $D_{i}$. A variable assignment $x_{i}=v$, $v\in D_{i}$, is treated as a token. A learnable embedding $\mathbf{e}(v)$ is assigned to each unique value $v\in\bigcup_{i=1}^{n}D_{i}$. The input variable assignment, represented as $\mathbf{X}=\{x_{1}=v_{1},x_{2}=v_{2},\dots,x_{n}=v_{n}\}$, forms the input token set to the Transformer. Thus, the input embedding set is given by
@@
Representing variable indices with Absolute Positional Encoding. Transformers use Absolute Positional Encoding (APE) to represent the position of tokens in a sequence. For CSPs, we use APE to encode the indices of variables. If the indices of a variable $x_{i}$ are multi-dimensional, we concatenate the positional encodings for each dimension. Specifically, let $x_{i_{1},i_{2},\dots,i_{k}}$ denote a variable with $k$-dimensional indices $(i_{1},i_{2},\dots,i_{k})$. The APE for this variable is computed as:
@@
where $\text{PE}(i_{k})$ is the positional encoding for dimension $k$. For example, in Sudoku, a variable $x_{12}$ would have an APE formed by concatenating the encodings for row 1 and column 2. This approach is inspired by the APE design in Vision Transformers (Carion et al., 2020; Li et al., 2025).
Constraint relations as Relative Positional Encoding. A Relative Positional Encoding (RPE) is typically used by Transformers to capture the positional relationship between tokens independently of their absolute positions in a sequence. For CSPs, we use RPE to encode the constraint relationships between variables. Specifically, we represent the CSP constraints as a binary constraint graph $\mathbf{G}=(V,E)$, where $V=\{1,2,\dots,n\}$ corresponds to the variables and $E$ contains edges between pairs of variables that participate together in at least one constraint of the CSP.
@@ 3.2 A Single-Step Transformer Architecture
where $\mathbf{e}(v_{i})$ is the token embedding of variable $x_{i}$’s current value $v_{i}$ as described in Equation 1, and $\text{APE}(x_{i})$ is its positional encoding as computed in Equation 2. $\alpha,\beta,\gamma$ are learnable scalars that allow the model to balance the contributions of each encoding, inspired by Li et al. (2025). The set of embeddings for all variables forms the input set $\mathbf{H}^{(0)}=\{\mathbf{h}_{1}^{(0)},\mathbf{h}_{2}^{(0)},\dots,\mathbf{h}_{n}^{(0)}\}$.
@@
ConsFormer employs a multi-head self-attention mechanism to compute updated representations of variables based on other variables. For each variable token $\mathbf{h}_{i}^{(l)}\in\mathbb{R}^{h\times 1}$ at layer $l$, the self-attention mechanism for a single attention head proceeds as follows:
@@ 3.3 Self-supervised Loss Function
where hyperparameter $\lambda_{i}$ is the weight assigned to $p_{i}$, and $f$ is an optional operation to transform the penalty for better learning. In practice, we implemented the common quadratic penalty, $f(x)=x^{2}$. The discrete constraints and their relaxed continuous counterparts we implemented for our experiments are shown in Table 1. Further discussion about the design process can be found in Appendix H and numerical examples of valid and invalid assignments for each constraint can be found in Appendix A.
@@ 3.4 Iterative Test-Time Deployment
In this sense, ConsFormer can be viewed as performing a single step of neural local search to improve the candidate solution. Our experiments focus on basic iterative solution refinement in one continuous sequence, without additional augmentations. However, this capability can be combined with techniques such as backtracking and random restarts (Rossi et al., 2006) to create a neuro-symbolic solver. Another possible extension is to incorporate ConsFormer as an evolutionary algorithm (Holland, 1992) utilizing the Transformer’s parallel processing ability to update a pool of candidate solutions all at once. While we leave these explorations as future work, we demonstrate the potential of this direction by implementing a simple multi-start strategy, which we show can significantly enhance performance in Appendix F.
@@ 4.1 Problems
Graph Coloring is the problem of finding an assignment of colors to vertices in a graph such that no two neighboring nodes share the same color. The problem is defined by the graph’s structure and the number of available colors $k$. We generate two sets of graph instances for $k=5$ and $k=10$ following a similar procedure as Tönshoff et al. (2023) (See Appendix C for details). Training graphs have 50 vertices for $k=5$ and 100 vertices for $k=10$ whereas OOD graphs have 100 for $k=5$ and 200 for $k=10$. We use inequality constraints of the form $x_{i}\neq x_{j}$ for an edge between nodes $i$ and $j$ and their penalty in Table 1 to represent the coloring constraints.
@@
MAXCUT aims to partition the vertices of a graph into two disjoint sets such that the number of edges crossing the partition is maximized. The problem can be viewed as a 2-coloring problem where the objective is to satisfy as many inequality constraints $x_{i}\neq x_{j}$ as possible. Following the same setup as Tönshoff et al. (2023), we generate random graphs with 100 vertices for training and evaluate generalization on benchmark instances from the GSET dataset (Ye, 2003), which includes weighted graphs with sizes ranging from 800 to 10000 vertices.
Table 2: Performance comparison for Sudoku. In-distribution test instances contain $1,000$ instances from the SATNet dataset, OOD refers to Out-of-Distribution evaluation on the RRN test dataset which contains 18K instances. *Values reported in (Du et al., 2024).
@@ 4.3 Results
Sudoku. Table 2 reports the performance of ConsFormer and various neural methods on the Sudoku task. ConsFormer solves 100% of the Sudoku tasks from the in-distribution test dataset. On the harder out-of-distribution dataset, ConsFormer significantly outperforms all learned methods, demonstrating superior generalization capabilities. ConsFormer achieves instance solve rates of 65.88% and 77.74% with 2k and 10k iterations, respectively.
@@
Graph Coloring. Table 3 summarizes the performance of OR-Tools, ANYCSP (Tönshoff et al., 2023), and ConsFormer on Graph Coloring instances. OR-Tools is a state-of-the-art traditional solver for constraint programming applications and serves as a strong baseline (Perron & Didier, ), achieving 100% on Sudoku instances. We ran test instances sequentially through all models with a 10-second timeout, though ConsFormer can process instances significantly faster if processed in batches due to the Transformer architecture. We note that the harder dataset is not out of distribution for OR-Tools, since it solves each task individually and is not a learning-based solver.
@@
On the more challenging Graph Coloring with $k=10$, ConsFormer surpasses OR-Tools in performance, showcasing the advantage of learned heuristic approaches: it may not surpass state-of-the-art solvers on smaller instances, but it excels in complex cases under short time limits—crucial for real-world applications. Surprisingly, ANYCSP failed to solve any instances within 10 seconds for both datasets, underscoring the scalability limitations of its graph-based representation and the difficulty of training with RL. Some additional baselines can be found in Appendix G.
Nurse Scheduling. For the Nurse Scheduling problem, ConsFormer matches OR-Tools in solving 100% of tasks across both in-distribution and out-of-distribution instances within the 10-second timeout. This high accuracy is expected, given the large number of feasible solutions for each instance, as detailed in Appendix C. However, solving this problem neurally is non-trivial, as the model must learn to balance multiple constraints within a single step of solution refinement. ANYCSP was not evaluated on this dataset due to their difficulty dealing with the AllDifferent constraint. This highlights ConsFormer’s potential to generalize to complex problems with diverse constraint structures.
MAXCUT. Table 4 compares various methods on GSET instances, reported as the relative gap (in percentage) to the best known cut values (Matsuda, 2019). ConsFormer and OR-Tools performances are obtained using the same set-up as ANYCSP, while the rest are computed directly from values reported by Tönshoff et al. (2023). While ANYCSP remains the best-performing method on GSET, ConsFormer achieves an average relative gap of 0.31% to 1.27% without extensive model tuning, demonstrating its ability to scale to larger problems with thousands of constraints.
Table 4: Performance comparison for MAXCUT tasks on GSET. Numbers reported are the average percentage gap to the best known cut size, the lower the better.
Method $|V|{=}800$ $|V|{=}1K$ $|V|{=}2K$ $|V|{\geq}3K$
Greedy 5.26 6.64 6.81 6.30
SDP 3.14 4.24 - -
RUNCSP 2.38 2.90 3.30 3.26
ECO-DQN 0.83 1.01 1.45 3.49
ECORD 0.11 0.16 0.36 1.53
ANYCSP 0.02 0.05 0.12 0.42
ConsFormer 0.31 0.34 0.43 1.27
OR-Tools 1.84 2.09 3.38 3.08
4.4 Ablations
Effect of subset improvement.